OpenStudy (anonymous):

How to find the slope of the curve f(x) = x^3 - 2 at x=1? I got m=1 but the answer is m=3. How do you get m=3? My work: m=(1+h)^3 - 2(1+h) - ((1)^3 - 2) / h = (h^3 + 3h^2 + 3h +1) - (2+2h) - (-1) / h = (h^3+3h^2+h) / h = h(h^2+3h+1) / h = (h^2+3h+1) = 1

OpenStudy (anonymous):

By taking derivative of it, gives you the \(\text{Slope}\) of the curve.. :)

OpenStudy (anonymous):

Derivative of Curve I mean.. :)

OpenStudy (anonymous):

Oh, I did not see that you have changed the question.. :P

OpenStudy (anonymous):

So, you have : \(f(x) = x^3 - 2\)

OpenStudy (anonymous):

Have you found its derivative?

OpenStudy (anonymous):

you can type your work here in the comments, I am looking down here to find your replies, and you are giving it in the question itself everytime.. :P

OpenStudy (anonymous):

I'm sorry, first time using website lol.

OpenStudy (anonymous):

I know, see just type your question above, and you can type your answer there too, but final answer.. Steps and discussion, we will do here in comments.. Okay?

OpenStudy (anonymous):

Now, don't change your question, I am seeing your steps.. :)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

See, there is no \(x\) with \(2\), right??

OpenStudy (anonymous):

Oh.. So it doesn't require the input

OpenStudy (anonymous):

yes, you should not multiply \((1+h)\) with \(2\), try without it..

OpenStudy (anonymous):

\[m=\frac{((1+h)^3 - 2) - ((1)^3 - 2)}{ h}\]

OpenStudy (anonymous):

See, you replace \(x\) by \((1+h)\), right?? so replace it, only where you are seeing \(x\)..

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

Yes, I understand

OpenStudy (anonymous):

Are you getting \(3\) now?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Wait.. I think I found my mistake

OpenStudy (anonymous):

Oh, go slow, the advantage is, you will commit less mistakes.. :) Don't worry, we learn from our mistakes only.. :)

OpenStudy (anonymous):

I got the answer!! :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[\huge \color{green}{\textsf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

Could you help me with another problem?

OpenStudy (anonymous):

you are welcome dear.. :)

OpenStudy (anonymous):

yeah sure..

OpenStudy (anonymous):

If you're not busy ;p

OpenStudy (anonymous):

Don't worry, if in case I am busy, then also you will keep getting help here, I assure you.. We also want to help those who are serious and come here for studying.. I found you serious as you showed your steps, told your answer, this means you have really tried this problem and unfortunately, you got some doubt, for which you are here.. :)

OpenStudy (anonymous):

I'm not sure if I'm suppose to multiply by the conjugate =/

OpenStudy (anonymous):

What is the question statement?

OpenStudy (anonymous):

\[f(x) = \sqrt{x-6}\]

OpenStudy (anonymous):

at x=14

OpenStudy (anonymous):

Whole statement?? Like your question what I think is : Find the value of \(f(x) = \sqrt{x-6}\) at \(x=14\) ?? Right?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Okay, are you seeing \(x\) there?? you are just replace that \(x\) with \(14\) as you are given \(x = 14\)..

OpenStudy (anonymous):

Oh dear, Electricity has gone, UPS is going to kill me and my pc now. @mayankdevnani please help here.. :)

OpenStudy (anonymous):

@mayankdevnani help here.. :)

OpenStudy (anonymous):

I did: m=\[\sqrt{14+h-6} - \sqrt{14-6}\] = \[\sqrt{h+8} - \sqrt{8} \]

OpenStudy (anonymous):

All divided by h

OpenStudy (mayankdevnani):

yeah !!! i am just reading the question

OpenStudy (mayankdevnani):

wait !!!

OpenStudy (mayankdevnani):

is this your question ? @thisissammy

OpenStudy (anonymous):

Yes

OpenStudy (mayankdevnani):

just plug the value of x=14 in the equation

OpenStudy (mayankdevnani):

and then simplify the equation!!

OpenStudy (anonymous):

*Find the slope of the curve

OpenStudy (mayankdevnani):

oh!!!

OpenStudy (mayankdevnani):

@thisissammy follow these steps :-

OpenStudy (mayankdevnani):

http://www.seriesmathstudy.com/tangentexample.htm please go through this site,this site must help you :) @thisissammy

OpenStudy (mayankdevnani):

@thisissammy first,find out the derivative of the given function

OpenStudy (mayankdevnani):

@waterineyes i do my best :)

OpenStudy (mayankdevnani):

are you there ? @thisissammy

OpenStudy (mayankdevnani):

http://www.seriesmathstudy.com/tangentexample.htm please go through this !!

OpenStudy (mayankdevnani):

first go through this ^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

Thank you for the sources

OpenStudy (mayankdevnani):

welcome :) i hope now you can find out derivative of any function

OpenStudy (mayankdevnani):

if not,then please take your time to learn and i hope now you can solve your question yourself :)

OpenStudy (anonymous):

Have you been taught the normal method for finding derivative? Or you have just learned finding the derivative using Fundamental Definition??

OpenStudy (anonymous):

I mean, I have seen you found the derivative using fundamental definition : using (1+h) and h all that..

OpenStudy (anonymous):

Other method is : \[\frac{d}{dx} x^n = n \cdot x^{n-1}\], although it is proved from fundamental definition.

OpenStudy (anonymous):

So: \(f(x) = x^3 - 2\), \(f'(x) = 3x^2 - 0\), derivative of constant is \(0\), so derivative of (-2) is \(0\)..

OpenStudy (anonymous):

So, \(f'(x) = 3x^2\), now put \(x= 1\), \(f'(1) = 3 = m\).. :)

OpenStudy (anonymous):

Getting ?

OpenStudy (anonymous):

That is what I was thinking that you asked first question on finding the slope of curve, then how could second question be so simple having just a value to plug in it.. :P

OpenStudy (anonymous):

So, I saw that you have reached to: \[m= \frac{\sqrt{h+8} - \sqrt{8}}{h}\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Of course with limit \(h\) tends to \(0\)..

OpenStudy (anonymous):

I haven't been taught the normal way yet

OpenStudy (anonymous):

But that is what I reached ^

OpenStudy (anonymous):

I am yet on that place only.. :P

OpenStudy (anonymous):

I have not used fundamental definition for derivatives, so I am not very good at it, I can do it other way, but you are yet to reach there.. Don't worry, if we will solve, then we will solve this only, no other way.. :)